An apple has 29 more calories than a peach , and 13 fewer calories than a banana. If three apples have 43 fewer calories than 2 bananas and 2 peaches , how many calories does each apple have?
Answer by stanbon(75887)

An apple has 29 more calories than a peach , and 13 fewer calories than a banana. If three apples have 43 fewer calories than 2 bananas and 2 peaches , how many calories does each apple have?
-----
Equations:
a = p + 29
a = b - 13
3a =(2b+2p)-43
-----
Rearrange:
a + 0 -p = 29
a - b +0 = -13
3a -2b -2p = -43
-----------------------
Use any method you know to get:
a = 75
b = 88
p = 46
==============
Cheers,
Stan H.